If magnitude of sum of two unit vector is 8/5.Then find magnitude of their difference?

Grade:12th pass

1 Answers

Aman gohel
6 years ago
Let unit vectors be A and BA+B=√(1+1+2cosX)[According to triangle`s law of vector addition]=>√(2+2cosX)=8/5=>CosX=7/25=>X=73.7°To find( A-B),X=90-73.7=16.3A-B=√(1+1+2cos16.3) =√(2+2cos16.3) =√4 =2(Ans)